James is 5 years older than Jennifer. Three years ago the product of their ages was 36 less than the product of their present ages. Find Jennifer's age.
Answer by ankor@dixie-net.com(22740)

Let a = James's present age
Let j = Jennifer's present age
:
Write an equation for each statement
:
James is 5 years older than Jennifer.
a = j + 5
:
Three years ago the product of their ages was 36 less than the product of their present ages.
(a-3)*(j-3) = a*j - 36
FOIL
aj - 3a - 3j + 9 = aj - 36
subtract aj from both sides
-3a - 3j = -36 - 9
mult by -1 get rid of all those negatives
3a + 3j = 45
replace a with (j+5) from the 1st statement
3(j+5) + 3j = 45
3j + 15 + 3j = 45
6j = 45 - 15
j = 30/6
j = 5 yrs is Jennifer's present age
:
:
See if that checks out in the statement James is 10 yrs old
"Three years ago the product of their ages was 36 less than the product of their present ages."
(10-3)(5-3) = 10*5 - 36
7 * 2 = 50 - 36: OK